# Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization

### v3

@@ 4.2 Optimization Performance
Ultimate Outcomes. We first examined the ultimate optimization outcomes of each method. The results are reported in Fig. 1 for synthetic benchmarks and in Fig. 2 for real-world benchmarks. Following [6], we used violin plots that illustrate the average of the obtained maximum function value over the ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region). We can observe that on average, the standard BO, including both SBO-1 and SBO-2, outperforms the competing methods at every target function, often by a large margin, except that in the cases of Hartman6(300, 6) and NAS201, the performance of SBO is slight worse than (but still very close to) the best baseline(s).
It is noteworthy that the performance of the competing high-dimensional BO methods appears to be sensitive to the structure within the target function. For example, SaasBO imposes a strong inductive bias that only a small number of variables is related to the objective (via the usage of the horse-shoe prior). The target function in Hartman6(300, 6) and DNA aligns well with this assumption. As a result, SaasBO exhibits outstanding performance in both tasks. However, when the objective deviates from this assumption, as seen in cases like Ackley(150, 150), StybTang(200, 200) and even Rosenbrock(300, 100), the performance of SaasBO experiences a dramatic drop. Another example is RDUCB and Tree-UCB, which assume that the target function can be decomposed as a summation of a group of functions, each over a small number of variables. RDUCB and Tree-UCB perform well in synthetic benchmarks that match this assumption, such as Rosenbrock(100, 100), Rosenbrock(300, 100), and StybTang(200, 200). However, their performance drops significantly in cases that do not align with their assumption, such as Ackley(150, 150) and Ackley(300, 150). Across all the benchmarks, the BNN based methods performed worse than most GP based methods. This might be due to that under low data regimes, it is more challenging for a neural network to capture the shape of the target function and quantify the uncertainty, especially for high-dimensional targets.
In contrast, the standard BO exhibits much greater robustness to the structure of the target function. Using standard GP equipped with Matérn kernels, both SBO-1 and SBO-2 consistently rank among the best in all tasks, regardless of whether the target function depends only on a subset of variables, can be decomposed, etc. Interestingly, as shown in Appendix Fig.7, even using a single length-scale parameter, the standard BO, i.e., SBO-3, still consistently shows superior performance, which closely matches ARD Matérn across most tasks, except for Ackley(300, 150), SBO-3 struggled in finding right locations to query, leading to poor performance.
@@
Runtime Performance. Next, we examined the runtime performance. For each method, we report the average maximum function value obtained at each step in Fig. 5 and Fig. 6 of the Appendix. We supplement the runtime standard deviation and the result of SBO-3 in Fig. 8 of the Appendix. Note that since we have twelve baseline methods, reporting the standard deviations of all the methods makes the figure very cluttered. To ensure clarity, in Fig. 8, we show the standard deviation of the best, worst, and median performed baselines so far at each step.
One can see that after a few initial steps, SBO-1 and SBO-2 (and SBO-3 in most cases) consistently generated superior queries, finding larger function values. Their curves are mostly above those of the competing methods, indicating a more efficient optimization of the target function. This trend is particularly evident in examples such as Rosenbrock(100, 100), StybTang(200, 200), Mopta08, Rover, SVM, and more.
The collective results demonstrate that the standard GP (with Matérn kernels) is effective for high-dimensional optimization. In comparison with more complex surrogate models using additional assumptions, the standard GP exhibits greater flexibility in adapting to the target function whose intrinsic structures are often agnostic in practice, and enables superior optimization results.
@@
4.3 Surrogate Learning
Next, we explicitly evaluated the performance of the standard GP in learning high-dimensional functions to confirm if it can serve as a reliable surrogate model for high-dimensional BO. To this end, from the optimization trajectories of all the methods, we randomly sampled 300 training examples from the six synthetic benchmark functions and four real-world benchmark tasks, including Mopta08 (124), Rover (60), DNA (180) and SVM (388). We sampled another 100 examples for testing. We compared with the surrogate models used in SaasBO, namely, placing a horse-shoe prior over the the length-scale parameters to prune the variables. We also compared with a random projection GP, denoted as ProjGP, to align with the baselines employing low-dimensional embeddings. ProjGP employs the kernel function in the form of $\vartheta({\bf x},{\bf x}^{\prime})=\kappa({\bf A}{\bf x},{\bf A}{\bf x}^{\prime})$, where ${\bf A}$ is a random projection matrix. We varied the rank of ${\bf A}$ from 10 and 20. Each element is sampled from the standard normal distribution. In addition, we compared with an additive GP, denoted as AddGP, where the kernel takes the form $\vartheta({\bf x},{\bf x}^{\prime})=\sum_{j}\kappa({\bf x}^{j},{\bf x}^{\prime j})$ where each ${\bf x}^{j}\subset{\bf x}$ and ${\bf x}^{\prime j}\subset{\bf x}^{\prime}$, ${\bf x}$ and ${\bf x}^{\prime}$ are the union of ${\bf x}^{j}$’s and ${\bf x}^{\prime j}$’s, respectively. AddGP imposes a decomposition structure in the functional space; see (5). We tested AddGP with randomly partitioning the input into 10 variables per group. We tested two versions: one using non-overlapping partitions, and the other overlapping partitions with one variable shared by successive groups. We trained the standard GP with ARD Matérn-5/2 kernel and with a single length-scale Matérn-5/2 kernel using MLE. SaasBO was trained by NUTS. For a fair comparison, ProjGP and AddGP were based on ARD Matérn-5/2 and trained via MLE as well.
(a) ARD Matérn
(b) Matérn
(c) SaasBO-NUTS
Figure 3: Alignment between the prediction (y-axis) and ground-truth function values (x-axis). The top row includes the results for Mopta08 (124) and the second row DNA (180).
We repeated the experiment for five times, and evaluated mean-square-error (MSE), the average test log-likelihood (LL), and the standard deviation. Table 4 and Appendix Table 5 report the results for all the GP models. The best and second best results are highlighted with a bold font. As we can see, in every dataset, standard GP with Mateŕ kernels achieves either the best or the second best result, demonstrating excellent and robust performance in learning high-dimensional functions.
In addition, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta08 (124) and DNA (180). As illustrated in Fig. 3, the prediction (and the predictive standard deviation) of ARD Matérn matches the ground-truth equally as well as SaasBO-NUTS, with a slight improvement over Matérn (using a single length-scale parameter).
Together these results demonstrate that directly incorporating high-dimensional inputs into the kernel function does not necessarily compromise the learning performance. This implies that, even when the input space is high dimensional, the popular Matérn kernel remains expressive enough to capture the similarities in the space, allowing them to perform well in covariance modeling. On the contrary, enforcing (unnecessary) low-rank structures or sparsity might cause information loss, potentially detrimental to optimization performance. Our observation is consistent with the recent work in kernel operator learning [3], where the standard kernel (GP) regression has demonstrated the state-of-the-art performance in learning function-to-function mappings. In this context, the input comprises of a discretized function, typically of very high-dimensionality (e.g., several thousands).
4.4 Computational Efficiency
So far, in our BO loop, the GP model is trained via MAP or MLE. Although the performance has been promising, it is worth investigating whether a full Bayesian approach would yield substantial improvements. In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction. We conducted another round of BO experiments. For each step, we used NUTS to train GP with ARD Matérn. We adopted the same settings as used in SaasBO-NUTS [6]: The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples. We show the running and final optimization performance in Fig 4 , and Appendix Fig. 9. It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is marginal. However, running NUTS is much more computationally costly compared to doing MAP and MLE. For example, on DNA, NUTS takes 59x and 40x running time of MAP and MLE, respectively, on average in each step. Therefore, for practical purposes, using MAP or MLE for GP training appears sufficient. It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
By contrast, for more complex surrogate models, such as SaasBO, the MCMC based training might be more crucial for achieving good performance. As depicted in Fig. 2, SaasBO-NUTS demonstrates a substantial improvement over SaasBO-MAP across all real-world benchmarks, excluding NAS201. However, the computational cost of running SaasBO-NUTS is considerably higher, ranging from 5.4x (NAS201) to 33.9x (SVM, DNA) compared to the running time of SaasBO-MAP. The substantial increase in computational costs may pose a bottleneck for practical usage. Detailed per-optimization-step running time for these methods are provided in Appendix Table 6 and 7.
Dataset ARD Matérn Matérn SaasBO Proj-GP Add-GP
Mopta08(124) 0.034$\pm$0.006 0.145$\pm$0.057 0.038$\pm$0.008 0.461$\pm$0.157 0.107$\pm$0.033
Rover(60) 0.327$\pm$0.078 0.284$\pm$0.09 0.284$\pm$0.09 0.548$\pm$0.097 0.321$\pm$0.097
DNA(180) 0.149$\pm$0.038 0.332$\pm$0.053 0.152$\pm$0.049 0.747$\pm$0.121 0.312$\pm$0.061
SVM(388) 0.034$\pm$0.01 0.103$\pm$0.017 0.034$\pm$0.014 0.392$\pm$0.071 0.077$\pm$0.012
Ackley(150, 150) 0.017$\pm$0.008 0.007$\pm$0.001 0.007$\pm$0.001 0.034$\pm$0.007 0.007$\pm$0.001
Ackley150(300, 150) 0.023$\pm$0.006 0.012$\pm$0.001 0.011$\pm$0.001 0.080$\pm$0.024 0.021$\pm$0.003
Hartmann6 (300, 6) 0.017$\pm$0.013 0.419$\pm$0.048 0.024$\pm$0.023 0.583$\pm$0.094 0.416$\pm$0.084
Rosenbrock(100, 100) 0.084$\pm$0.013 0.118$\pm$0.023 0.124$\pm$0.025 0.378$\pm$0.069 0.097$\pm$0.015
Rosenbrock(300, 100) 0.090$\pm$0.064 0.160$\pm$0.09 0.131$\pm$0.075 0.533$\pm$0.214 0.138$\pm$0.075
Stybtang(200, 200) 0.043$\pm$0.018 0.036$\pm$0.014 0.039$\pm$0.012 0.253$\pm$0.040 0.038$\pm$0.009
Table 4: Mean Squared Error(MSE) of each surrogate model. The results were averaged from five runs.
Figure 4: Optimization performance when using ARD Matérn trained with MAP and MLE, and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
4.5 Ablation Study and Failure Modes with Other Settings
Finally, we conducted a careful ablation study on the alternative settings to carry out standard BO. We found, however, those settings, such as using SE kernel and EI acquisition, often lead to inferior performance in high-dimensional optimization, which align with the a few failure cases reported in the literature [15, 20]. The detailed results and analysis are given in Section B of the Appendix. We summarize the favorable and unfavorable settings for high-dimensional optimization in Table 1.
@@ 5 Conclusion
We have presented a comprehensive investigation on standard BO (SBO) in high-dimensional optimization. We discovered that simply using Matérn kernel and UCB, SBO has already achieved the state-of-the-art performance. There seems no need for imposing additional structure assumptions for function learning and/or placing strong priors over the model parameters. In the future, we will continue applying SBO in more tasks, and seek for theoretical justifications and guarantees.
Appendix A More Experiment Details
A.1 Definition of Synthetic Functions
@@
Figure 5: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in synthetic benchmarks. The results were averaged from 10 runs.
Figure 6: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in real-world benchmarks. The results were averaged from 10 runs.
Figure 7: Ultimate optimization performance: Maximum Function Value obtained at the end of optimization loop, including SBO-3, namely training with a single length-scale parameter in Matérn-5/2 kernel. The results were collect from 10 runs.
Figure 8: Runtime performance: Maximum Function Value obtained vs. Number of Steps in real-world benchmarks, including SBO-3, namely training with a single length-scale parameter in Matérn-5/2 kernel. The results were averaged from 10 runs. The shaded region depicts the standard deviation.
A.2 Implementation
@@ A.3 Computing Environment
Task ARD Matérn Matérn SaasBO Proj-GP Add-GP
Mopta08(124) 0.570$\pm$0.08 -0.277$\pm$0.101 0.560$\pm$0.064 -0.962$\pm$0.183 -0.089$\pm$0.105
Rover(60) -0.975$\pm$0.216 -0.647$\pm$0.158 -0.638$\pm$0.14 -0.995$\pm$0.116 -0.703$\pm$0.125
DNA(180) -0.256$\pm$0.261 -0.666$\pm$0.102 -0.183$\pm$0.158 -1.137$\pm$0.143 -0.679$\pm$0.143
SVM(388) 0.581$\pm$0.119 -0.179$\pm$0.042 0.650$\pm$0.12 -0.837$\pm$0.095 0.062$\pm$0.129
Ackley(150, 150) 0.428$\pm$0.222 0.633$\pm$0.067 0.574$\pm$0.068 -0.006$\pm$0.268 0.569$\pm$0.066
Ackley(300, 150) 0.232$\pm$0.155 0.252$\pm$0.057 0.279$\pm$0.077 -0.372$\pm$0.185 0.220$\pm$0.074
Hartmann6(300, 6) 1.200$\pm$0.169 -0.625$\pm$0.079 1.528$\pm$0.564 -0.927$\pm$0.084 -0.617$\pm$0.118
Rosenbrock(100, 100) 0.640$\pm$0.081 0.445$\pm$0.062 0.526$\pm$0.08 -0.466$\pm$0.182 0.521$\pm$0.057
Rosenbrock(300, 100) 0.518$\pm$0.225 0.121$\pm$0.155 0.357$\pm$0.21 -0.706$\pm$0.209 0.189$\pm$0.2
Stybtang(200, 200) 0.574$\pm$0.164 0.563$\pm$0.073 0.543$\pm$0.067 -0.458$\pm$0.069 0.582$\pm$0.072
Table 5: Test log likelihood of each surrogate model. The results were averaged over five runs. Bolded are the best and second best results.
@@
Figure 9: Optimization performance using MLE training, MAP training with a diffuse prior over the length-scale parameters, and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
Appendix B Ablation Study and Failure Modes With Other Settings
In this section, we conducted a careful ablation study to investigate alternative popular settings in executing standard BO. From these settings, we identified a few failure modes during high-dimensional optimization. We provided analysis and discussions accordingly.
B.1 GP Training
First, we examined MLE training with LBFGS. We found that, however, LBFGS often incurs numerical issues, and cannot reliably finish model training. This is consistent with the report in other works [12]. The issue might be because that without constraints (or regularization) over the parameters (such as the length-scales and amplitude), the exploration of LBFGS tends to be overly aggressive, which easily cause numerical problems. Hence, using stochastic optimization, like ADAM, might be a wise choice for MLE training.

### v4

@@ 6.2 Optimization Performance
Ultimate Outcomes. We first examined the ultimate optimization outcomes of each method. The results are reported in Fig. 3 for synthetic benchmarks and in Fig. 4 for real-world benchmarks. Due to space limit, we supplement the BNN-based BO results in Appendix Section C.1. Following [6], we used violin plots that illustrate the average of the obtained maximum function value over the ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region).
On average, SBO-Matérn, SBO-SE(RI) and VBO consistently deliver top-tier results, often outperforming the other methods by a large margin, e.g., on benchmarks Ackley (300, 150), DNA, and SVM. It is noteworthy that SBO-SE also performed exceptionally well when the input dimension $d\leq 180$, as seen in benchmarks Rosenbrock (100, 100), Mopta08, NAS, and DNA. However, as $d$ increases, e.g., Ackley (300, 150), Hartman6 (300, 6), Rosenbrock (300, 100), StybTang(200, 200) and SVM (388), the performance of SBO-SE dropped dramatically due to the gradient vanishing issue (see our theoretical analysis and numerical experiments in Section 4). After applying our robust initialization, the performance of SE kernels, namely, SBO-SE (RI) immediately recovered to top-tier levels, closely matching both SBO-Matérn and VBO. On the other hand, though SBO-Matérn used a small length-scale initialization ($\ell_{0}=0.693$), the Matérn kernel’s structure has made it far less susceptible to the gradient vanishing issue, enabling it to handle much higher dimensions than the SE kernel. Free from this issue, SBO-Matérn consistently achieved top-tier results.
@@
Runtime Performance. Next, we examined the runtime performance. Fig. 5 reports the average maximum function value obtained at each step and the standard deviation across all runs. Since we have 16 methods for comparison, showing the standard deviation for every method makes the figure cluttered. To ensure clarify, we show the curves and standard deviation of SBO, VBO, and the best, worst, and median performed remaining baselines so far at each step.
After an initial stage, SBO-Matérn, SBO-SE (RI), and VBO consistently produced superior queries, achieving larger function values. Their performance curves are generally above those of competing methods, reflecting a more efficient optimization of the target function. This trend is particularly noticeable in benchmarks such as Ackley (300, 150), Rosenbrock (300, 100), and SVM. When $d>180$, however, the performance of SBO-SE declines, with its curve positioned toward the bottom due to the gradient vanishing issue. In benchmarks such as Rosenbrock (100, 100) and DNA (180), the curve for VBO is predominantly above those of SBO-Matérn and SBO-SE (RI), indicating even more efficient optimization. However, on Ackley (150, 150) and Ackley (300, 150), VBO’s curve is mostly below that of SBO-Matérn and SBO-SE (RI). For the remaining benchmarks, their curves are close and often overlap. Overall, in all cases, they tend to converge to similar function values.
These results collectively demonstrate that standard BO can excel in high-dimensional optimization problems. The lack of additional structural assumptions in GP modeling may make it more flexible in capturing various correlations within high-dimensional spaces. The primary challenge for standard BO arises from the gradient vanishing issue, which is often due to improper initialization of the length-scales. By using our robust initialization method, this risk can be significantly mitigated.
Additional Results. We conducted extensive additional evaluations and ablation studies to examine the impact of alternative GP training methods, exploration parameters in UCB, and other acquisition functions on SBO performance in high-dimensional settings. We found that MLE achieves optimization results comparable to the MAP training employed in our evaluation. However, using concentrated priors on the length-scale parameters (concentrated on small mean values) — such as the popular Gamma(3,6) prior that is the default choice in BOTorch — often leads to training failure and poor performance, aligning with our theoretical analysis. Both MLE and MAP perform similarly to Markov Chain Monte Carlo (MCMC) training in terms of optimization outcomes, but MCMC is significantly slower. UCB is notably robust to variations in the exploration parameter. Additionally, log-EI facilitates strong performance in high-dimensional optimization, whereas EI and TS often lead to subpar results. Detailed results and discussions are provided in Appendix Section C.
@@ 7 Conclusion
We conducted a thorough investigation of standard BO in high-dimensional optimization, both empirically and theoretically. Our analysis identified gradient vanishing as a major failure mode of standard BO, and we proposed a simple yet robust initialization method to address this issue. Our empirical evaluation demonstrates that, once this problem is mitigated, standard BO can achieve state-of-the-art performance in high-dimensional optimization.
Appendix A Proofs
A.1 Proof of Proposition 4.1
Proof.
We intend to solve $\rho^{2}/e^{\rho^{2}}<\xi$, which is equivalent to solving
$\log r-r<\log\xi,$
where $r=\rho^{2}$. We then leverage the fact that $\log r<\sqrt{r}$ when $r>1$. Since $1/e>\xi$, we know the solution of $\rho$ must be larger than 1, implying that the solution for $r$ is larger than $1$. To meet the inequality, we can solve
$\log r-r<\sqrt{r}-r<\log\xi.$
It then converts to a quadratic form,
$\sqrt{r}-\sqrt{r}^{2}<\log\xi.$ (10)
Note that $\rho=\sqrt{r}$. It is straightforward to solve this inequality and we can obtain that the inequality holds when
$\rho>\tau_{\text{SE}}=\frac{1}{2}+\sqrt{\frac{1}{4}-\log\xi}.$
A.2 Proof of Lemma 4.2
Proof.
First, since each $x_{ik}$ and $x_{jk}$ independently follow $\text{Uniform}(0,1)$, it is straightforward to obtain $\mathbb{E}[(x_{ik}-x_{jk})^{2}]=\frac{1}{6}$. Let us define $\gamma=\|{\bf x}_{i}-{\bf x}_{j}\|^{2}$. Then we have $\mathbb{E}[\gamma]=\frac{1}{6}d$. According to Hoeffding’s inequality, for all $t>0$,
$p(|\gamma-\frac{1}{6}d|\geq t)\leq 2\exp\left(-\frac{2t^{2}}{d}\right).$ (11)
Since $\rho=\frac{\sqrt{\gamma}}{\ell_{0}}$, we have $p(\rho>\tau)=p(\gamma>\tau^{2}\ell_{0}^{2})$. Since $d>6\ell_{0}^{2}\tau^{2}$, we set $t=\frac{1}{6}d-\ell_{0}^{2}\tau^{2}>0$, and apply (11) to obtain
$p(|\gamma-\frac{1}{6}d|\geq\frac{1}{6}d-\ell_{0}^{2}\tau^{2}|)\leq 2\exp\left(-\frac{2\left(\frac{1}{6}d-\ell_{0}^{2}\tau^{2}\right)^{2}}{d}\right)=2\exp\left(-\frac{\left(d-6\ell_{0}^{2}\tau^{2}\right)^{2}}{18d}\right),$
which is equivalent to
$p(|\gamma-\frac{1}{6}d|<\frac{1}{6}d-\ell_{0}^{2}\tau^{2})>1-2\exp\left(-\frac{\left(d-6\ell_{0}^{2}\tau^{2}\right)^{2}}{18d}\right).$ (12)
Since $p(\gamma>\ell_{0}^{2}\tau^{2})\geq p(\ell_{0}^{2}\tau^{2}<\gamma<\frac{1}{3}d-\ell_{0}^{2}\tau^{2})=p(|\gamma-\frac{1}{6}d|<\frac{1}{6}d-\ell_{0}^{2}\tau^{2})$, combining with (12) we obtain the bound
$p(\rho>\tau)=p(\gamma>\ell_{0}^{2}\tau^{2})>1-2\exp\left(-\frac{\left(d-6\ell_{0}^{2}\tau^{2}\right)^{2}}{18d}\right).$
A.3 Proof of Proposition 4.3
Proof.
To solve $\rho^{2}/e^{\sqrt{5}\rho}<\xi$, it is equivalent to solving
$\frac{1}{5}\frac{r^{2}}{e^{r}}<\xi,\;\;\;r=\sqrt{5}\rho,$
which is further equivalent to
$\log\frac{1}{5}+2\log r-r<\log\xi.$
We leverage $\log r<\sqrt{r}$ for $r>1$, and solve the upper bound,
$\log\frac{1}{5}+2\sqrt{r}-r<\log\xi,$
which gives
[…]
A.4 Proof of Lemma 5.1
Proof.
First, similar to the proof in Section A.2, under the uniform distribution on $[0,1]^{d}$, from Hoeffding’s inequality, we have for all $t>0$,
$p(|\gamma-\frac{1}{6}d|\geq t)\leq 2\exp\left(-\frac{2t^{2}}{d}\right).$ (14)
When we set $\ell_{0}=c\sqrt{d}$ such that $c>1/\sqrt{6}\tau$, we have $\ell_{0}^{2}\tau^{2}>d/6$. Let us choose $t=\ell_{0}^{2}\tau^{2}-d/6=c^{2}\tau^{2}d-d/6>0$, and apply the Hoeffding’s inequality (14),
$p\left(|\gamma-d/6|\geq\ell_{0}^{2}\tau^{2}-d/6\right)\leq 2\exp\left(-\frac{2(\ell_{0}^{2}\tau^{2}-d/6)^{2}}{d}\right)=2\exp\left(-2(c^{2}\tau^{2}-\frac{1}{6})^{2}d\right).$ (15)
Since $p(\rho\geq\tau)=p(\gamma\geq\ell_{0}^{2}\tau^{2})\leq p(|\gamma-d/6|\geq\ell_{0}^{2}\tau^{2}-d/6)$, we obtain
$p(\rho\geq\tau)\leq 2\exp\left(-2(c^{2}\tau^{2}-\frac{1}{6})^{2}d\right).$
Appendix B More Experiment Details
B.1 Definition of Synthetic Functions
@@
Figure 5: Runtime performance: Maximum Function Value Obtained vs. Number of Steps in all real-world and synthetic benchmarks. The results were averaged from 10 runs. For cleaner view, we plotted the full results only for SBO and VBO. For the remaining methods, we show best, median, and worst result at each step.
Figure 6: Runtime performance comparison with BNN-based BO methods and SBO with a single length-scale parameter. Matérn kernel was used to conduct SBO. For clarity, VBO was excluded for comparison. The results were averaged from 10 runs. The shaded region depicts the standard deviation.
B.2 Implementation
@@ B.3 Computing Environment
Method Structural Assumption
ALEBO, HESBO Low-dim embedding
SaasBO Partial variable dependency
RDUCB, Tree UCB Additive function decomposition
SBO, TURBO, BNN None
@@
Figure 7: Optimization performance using MLE training, MAP training with a uniform prior over the length-scale parameters, and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
Figure 8: Optimization performance using MLE training, MAP training with a uniform prior over the length-scale parameters, and NUTS on the rest of benchmarks. The results were summarized from 10 runs.
Appendix C Additional Results
C.1 BNN-based BO and Standard BO with a Single Length-Scale
Due to the space limit, we did not report the results of the BNN-based BO in the main paper. Here, we supplement the comparison with the aforementioned two BNN-based BO approaches, BNN Laplace and BNN HMC, as specified in Section B.2. In addition, we also tested SBO with a single length-scale. We used the Matérn kernel. The results are reported in Fig. 6. Across all the benchmarks, the performance of the BNN based methods is worse than the median of GP based methods. In cases, such as SVM, Hartmann6 (300, 6) and Rosenbrock, the performance of the BNN based methods is close to or overlapping with the performance of the worst GP baselines. This might be due to that under low data regimes, it is more challenging for a neural network to capture the shape of the target function and quantify the uncertainty, especially for high-dimensional targets.
Interestingly, even using a single length-scale parameter, the standard BO, namely SBO-Single — as long as it does not encounter gradient vanishing — still consistently shows superior performance, which closely matches ARD Matérn across most tasks, except for Ackley(300, 150), SBO-Single struggled in finding right locations to query, leading to poor performance.
C.2 Alternative Ways to GP Training
Figure 9: Final optimization performance using a diffuse prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
Figure 10: Runtime optimization performance using a uniform prior, Uniform(0.001, 30), vs. a concentrated Gamma prior, Gamma(3,6), over the length scale parameters.
C.2.1 Maximum Likelihood Estimation (MLE)
In the main paper, we report the SBO performance when training GP via MAP estimation. We then examined another popular GP training approach, MLE training. We used ADAM optimization [17] [^2] We also experimented with using L-BFGS for MLE training. However, L-BFGS frequently encounters numerical issues and often fails to reliably complete model training. This is consistent with findings reported in [13]. The issue likely arises because, without constraints or regularization on parameters like the length-scales and amplitude, L-BFGS tends to explore too aggressively, leading to numerical instability. to perform MLE, with learning rate $0.1$ and the maximum number of epochs was set to $500$. We employed ARD Matérn kernel. The ultimate and running optimization performance is shown in Fig.  9 and 10. As we can see, the BO performance of using MAP and MLE training is close, and in many cases significantly overlaps, such as in DNA, SVM, Ackley (150, 150) and StybTang (200, 200).
C.2.2 MAP Training with Alternative Priors